# Effect of generic dark matter halo on transonic accretion onto galactic black holes

### v1

@@ VII Radiative emission properties
On the contrary, for fixed compactness, the bolometric luminosity increases with the halo mass, but it soon saturates with the NFW1 profile, giving the maximum luminosity followed by the Hernquist, NFW, Einasto and HQ-type DM spike distributions. The variation of the bolometric luminosity for the HQ-type DM spike distribution with the halo mass shows a rather curious behaviour when compared to other profiles (see the right panel of Fig. 9). For sufficiently high halo mass, the bolometric luminosity of the HQ-type DM spike distribution is less than that of a vacuum Schwarzschild BH of the same mass. This initial variation of the bolometric luminosity with the halo mass for a HQ-type DM spike distribution of constant compactness $(\sim 0.01)$ is related to the variation of the halo density profile and the associated change in geometry and properties of the accretion flow with the halo mass as discussed in Sections VI and B. At a high enough halo mass, the relative disk height attains unity at a much smaller radius, reducing the luminosity even below the Schwarzschild limit.
@@
In Fig. 10, we plot the maximum luminosity of the galactic BH with different DM profiles for a constant halo mass of $100M_{\rm BH}$ and halo compactness $0.01$ with $M_{\rm BH}=5\times 10^{8}~{}M_{\odot}$. In the plot, we vary the specific flow energy while maintaining the specific angular momentum of the flow slightly less than the Keplerian angular momentum in each case. We observe that the maximum bolometric luminosity is obtained for the NFW1 distribution, followed by the Einasto, Hernquist and NFW distributions. The maximum luminosity in each of these cases is higher than the maximum bolometric luminosity of a vacuum Schwarzschild BH. However, the bolometric luminosity of the HQ-type DM spike distribution around an equal mass BH with the same set of halo parameters show a significantly lower luminosity. This low luminosity is a manifestation of the fact that at the chosen compactness and halo mass, the critical radius $r_{0}$ is very small (see Fig. 7).
@@ VIII Summary and Discussion
In this work, we study the advection dominated transonic hot accretion flow of an inviscid relativistic fluid onto a galactic SMBH described by the metric (Eq. 1). We model the geometry of the galactic BH as a spherically symmetric central Schwarzschild BH immersed in a DM halo with radial density distributions. We represent the DM halo as an ideal gas composed of cold collision-less DM particles. In the absence of hierarchical mergers Merritt et al. (2002), our assumptions imply a seed BH at the centre of a galaxy, dressed with the relativistic distribution of DM Sadeghian et al. (2013) and a density spike in the inner region Vasiliev (2007); Shapiro and Shelton (2016); Ullio et al. (2001); Merritt (2004); Gnedin and Primack (2004); Bertone and Merritt (2005). Following Speeney et al. (2024), we have modelled this DM density spike in two different ways: (i) Hernquist, NFW, and Einasto distributions, truncated at $r=4M_{\rm BH}$ (ensuring that all energy conditions are obeyed) and (ii) a fully relativistic HQ-type DM spike model where we assume the HQ-type DM spike grows adiabatically from an initial Hernquist type distribution. We explicitly assume the central BH to grow adiabatically, with the accretion rate of the DM particles being negligible (Fig. 1). This leads to the anisotropic nature (zero radial pressure) of the DM distributions, a key ingredient in the entire analysis.
@@
Although the presence of the DM halo does not change the location of the event horizon ($r_{\rm h}=2M_{\rm BH}$), it greatly modifies the geometry of the surrounding spacetime (including the location of the lightrings and ISCO). This change in geometry, in turn, reduces the effective gravitational potential experienced by an infalling fluid element (Fig. 2). The Keplerian angular momentum also shows a significant deviation from that of Schwarzschild BH, allowing flow with higher specific angular momenta to reach the event horizon (Fig. 11). We observe that in all the cases, the presence of DM halo shifts the critical points to smaller radii, affecting the solution topologies (Figs. 5, 6 and 4). The primary observable in our study is the luminosity of the emitted Bremsstrahlung radiation. Since the bolometric luminosity for I-type accretion flow is known to be the highest Patra et al. (2024); Sen et al. (2024), we chose the flow parameters such that the flow passes only through the inner saddle type critical point.
@@
We observe (Fig. 7) that the gravitational effect of the DM halo weakens the inflow of the accreting fluid resulting in reduced flow velocity. The gravitational effect of the DM halo also increases the local flow temperature compared to that of a Schwarzschild BH. Since the mass accretion rate is assumed to remain constant throughout the accretion flow, the local mass density of the accreting fluid is always inversely proportional to the local velocity of the flow. Again, since the specific energy of the flow remains constant, the lower flow velocity implies a lower kinetic energy and higher thermal energy of the accreting fluid, resulting in an increased local temperature (particularly, close to the inner edge of the disk). The increase in local flow temperature results in an increased thermal pressure, which increases the relative disk height, making the accretion disk quasi-spherical. Though the optical depth ($\tau$) of the disk still remains small, the increased disk height reduces the effective portion of the disk, contributing to the disk luminosity. This is more prominent for highly compact halos with a higher halo mass. The SED also shows a significant deviation from that of the vacuum Schwarzschild BH (Fig. 8). Since the dominant contribution to the luminosity for I-type accretion comes from the inner portion of the accretion disk, where deviation from the vacuum Schwarzschild geometry is strongly governed by the presence and nature of DM spike (overdensity), the variation in luminosity for the same set of halo and flow parameters indeed encode information about the nature of the density spike. The variations of the luminosity of the emitted radiation and properties of the accretion flow increase with the mass and compactness of the DM halo (Fig. 9). However, in the case of the HQ-type DM spike distribution, the nonlinear dependence of the density profile on the halo mass changes the flow properties significantly with the halo mass (see Appendix B for details).
@@
Indeed, the presence of DM and the nature of its distribution around the SMBH strongly influences the geometry of the surrounding spacetime. However, the variation of bolometric luminosity may not be significant to infer the exact nature of the DM spike close to the BH. Thus, further research is required on more accurate modelling of the DM density spike around the SMBH and its correlation with a more realistic accretion model. In this regard, we must also emphasize the limitations of our accretion model. As a first step to address this complicated problem involving multiple length and time scales, we assume the accretion flow to be governed by ideal fluid dynamics, free of any dissipation mechanism. This results in a constant angular momentum of the fluid throughout the accretion flow. Incorporating dissipation mechanisms, such as radiative cooling of the disk and thermal conduction, will result in regulating the disk temperature and, hence, the disk height influencing the spectral emission properties. Furthermore, throughout our analysis, we assume the electron and ion temperatures to be related by a simple scaling factor. We also assume the absence of long-range magnetic field.
The present work can be extended in several ways by systematically relaxing the assumptions as discussed above. The most interesting will be to see the effect of rotation of the central BH with the presence of a long-range magnetic field, cooling mechanism and the effect of the DM distribution on the shock solution. However, such an analysis is beyond the scope of the present work and will be communicated separately. Finally, since the gravitational interaction of DM is the same as that of baryonic matter, one must also argue the validity of the assumption of negligible accretion rate of the DM particle as compared to that of baryonic matter.
@@ IX Acknowledgment
The work of AC is supported by the National Postdoctoral Fellowship of the Science and Engineering Research Board (SERB), ANRF, Govt. of India (File No.: PDF/2023/000550). SC acknowledges support of MATRICS research grant awarded by Science and Engineering Research Board (SERB), ANRF, Govt. of India through grant no. MTR/2022/000318.
@@ Appendix A Keplerian angular momentum
This implies that matter rotating in a Keplerian orbit will stay in equilibrium as the centrifugal force balances the effective gravitational attraction. The variation of the Keplerian angular momentum for the different DM profiles is shown in Fig. 11. We note that the presence of the DM halo always increases the Keplerian angular momentum from that of Schwarzschild BH. This implies that advection dominated transonic hot accretion flow in the presence of DM halo can have higher specific angular momentum. Except for highly compact halos, the Einasto distribution shows the highest Keplerian angular momentum at any radius.
The minima of the Keplerian angular momenta correspond to the radius of the marginally stable orbit Abramowicz and Fragile (2013); Rezzolla and Zanotti (2013). We observe that the radius of the marginally stable orbit (ISCO, see Eq. 7) shifts inwards due to the presence of the DM halo signalling a weaker inflow. This is consistent with the observation in Heydari-Fard et al. (2025).

### v2

@@ VII Radiative emission properties
On the contrary, for fixed compactness, the bolometric luminosity increases with the halo mass, but it soon saturates with the NFW1 profile, giving the maximum luminosity followed by the Hernquist, NFW, Einasto and HQ-type DM spike distributions. The variation of the bolometric luminosity for the HQ-type DM spike distribution with the halo mass shows a rather curious behaviour when compared to other profiles (see the right panel of Fig.˜9). For sufficiently high halo mass, the bolometric luminosity of the HQ-type DM spike distribution is less than that of a vacuum Schwarzschild BH of the same mass. This initial variation of the bolometric luminosity with the halo mass for a HQ-type DM spike distribution of constant compactness $(\sim 0.01)$ is related to the variation of the halo density profile and the associated change in geometry and properties of the accretion flow with the halo mass as discussed in Sections˜VI and B. At a high enough halo mass, the relative disk height attains unity at a much smaller radius, reducing the luminosity even below the Schwarzschild limit.
@@
In Fig.˜10, we plot the maximum luminosity of the galactic BH with different DM profiles for a constant halo mass of $100M_{\rm BH}$ and halo compactness $0.01$ with $M_{\rm BH}=5\times 10^{8}~M_{\odot}$. In the plot, we vary the specific flow energy while maintaining the specific angular momentum of the flow slightly less than the Keplerian angular momentum in each case. We observe that the maximum bolometric luminosity is obtained for the NFW1 distribution, followed by the Einasto, Hernquist and NFW distributions. The maximum luminosity in each of these cases is higher than the maximum bolometric luminosity of a vacuum Schwarzschild BH. However, the bolometric luminosity of the HQ-type DM spike distribution around an equal mass BH with the same set of halo parameters shows a significantly lower luminosity. This low luminosity is a manifestation of the fact that at the chosen compactness and halo mass, the critical radius $r_{0}$ is very small (see Fig.˜7).
@@ VIII Summary and Discussion
In this work, we study the advection-dominated transonic hot accretion flow of an inviscid relativistic fluid onto a galactic SMBH described by the metric (Eq.˜1). We model the geometry of the galactic BH as a spherically symmetric central Schwarzschild BH immersed in a DM halo with radial density distributions. We represent the DM halo as an ideal gas composed of cold collisionless DM particles. In the absence of hierarchical mergers Merritt et al. (2002), our assumptions imply a seed BH at the centre of a galaxy, dressed with the relativistic distribution of DM Sadeghian et al. (2013) and a density spike in the inner region Vasiliev (2007); Shapiro and Shelton (2016); Ullio et al. (2001); Merritt (2004); Gnedin and Primack (2004); Bertone and Merritt (2005). Following Speeney et al. (2024), we have modelled this DM density spike in two different ways: (i) Hernquist, NFW, and Einasto distributions, truncated at $r=4M_{\rm BH}$ and (ii) a fully relativistic HQ-type DM spike model where we assume the HQ-type DM spike grows adiabatically from an initial Hernquist type distribution. We explicitly assume the central BH to grow adiabatically (Fig.˜1). To obtain the effective spacetime metric in the presence of DM, we use the Einstein cluster formalism. This leads to the anisotropic nature (zero radial pressure) of the DM Einstein (1939); Geralico et al. (2012). Thus, within the timescale of baryonic accretion, our analysis assumes the accretion rate of the DM particles being negligible. Furthermore, the analysis also ensures that the stress-energy tensor remains covariantly conserved and all the energy conditions are satisfied.
@@
Although the presence of the DM halo does not change the location of the event horizon ($r_{\rm h}=2M_{\rm BH}$), it greatly modifies the geometry of the surrounding spacetime (including the location of the lightrings and ISCO). This change in geometry, in turn, reduces the effective gravitational potential experienced by an infalling fluid element (Fig.˜2). The Keplerian angular momentum also shows a significant deviation from that of Schwarzschild BH, allowing flow with higher specific angular momenta to reach the event horizon (Fig.˜11). We observe that in all the cases, the presence of DM halo shifts the critical points to smaller radii, affecting the solution topologies (Figs.˜5, 6 and 4). The primary observable in our study is the luminosity of the emitted Bremsstrahlung radiation. Since the bolometric luminosity for I-type accretion flow is known to be the highest Patra et al. (2024); Sen et al. (2024), we chose the flow parameters such that the flow passes only through the inner saddle type critical point.
@@
We observe (Fig.˜7) that the gravitational effect of the DM halo weakens the inflow of the accreting fluid, resulting in reduced flow velocity. The gravitational effect of the DM halo also increases the local flow temperature compared to that of a Schwarzschild BH. Since the mass accretion rate is assumed to remain constant throughout the accretion flow, the local mass density of the accreting fluid is always inversely proportional to the local velocity of the flow. Again, since the specific energy of the flow remains constant, the lower flow velocity implies a lower kinetic energy and higher thermal energy of the accreting fluid, resulting in an increased local temperature (particularly, close to the inner edge of the disk). The increase in local flow temperature results in an increased thermal pressure, which increases the relative disk height, making the accretion disk quasi-spherical. Though the optical depth ($\tau$) of the disk still remains small, the increased disk height reduces the effective portion of the disk, contributing to the disk luminosity. This is more prominent for highly compact halos with a higher halo mass. The SED also shows a significant deviation from that of the vacuum Schwarzschild BH (Fig.˜8). Since the dominant contribution to the luminosity for I-type accretion comes from the inner portion of the accretion disk, where deviation from the vacuum Schwarzschild geometry is strongly governed by the presence and nature of DM spike (overdensity), the variation in luminosity for the same set of halo and flow parameters indeed encode information about the nature of the density spike. The variations of the luminosity of the emitted radiation and properties of the accretion flow increase with the mass and compactness of the DM halo (Fig.˜9). However, in the case of the HQ-type DM spike distribution, the nonlinear dependence of the density profile on the halo mass changes the flow properties significantly with the halo mass (see Appendix˜B for details).
@@
Indeed, the presence of DM and the nature of its distribution around the SMBH strongly influence the geometry of the surrounding spacetime. However, the variation of bolometric luminosity may not be significant to infer the exact nature of the DM spike close to the BH. Thus, further research is required on more accurate modelling of the DM density spike around the SMBH and its correlation with a more realistic accretion model. In this regard, we mention the limitations of the accretion model used in our work. As a first step to address the present problem involving multiple length and time scales, we assume the accretion flow to be governed by ideal fluid dynamics, free from any dissipation mechanism. This results in a constant angular momentum of the fluid throughout the accretion flow as viscosity is ignored. Note that in the inner regions of the disc, the viscous timescale often exceeds the infall timescale of the accreting matter. As a result, the flow does not have sufficient time to redistribute angular momentum outward through viscous interactions, rendering the inner accretion effectively inviscid Fukue (1987); Chakrabarti (1989)(see also Appendix A of Dihingia et al. (2019)).
Incorporating other dissipation mechanisms, such as radiative cooling and thermal conduction, are expected to regulate the disk temperature, influencing the spectral emission properties. Furthermore, throughout our analysis, we consider the electron and ion temperatures to be related by a simple scaling factor. We also disregard the presence of structured magnetic fields in this work.
The present findings obtained using these considerations are expected to be in agreement at least qualitatively with the observationally inferred systems such as Sgr A* or M87*, where magnetic fields, turbulance, viscosity, and multi-temperature plasma dynamics play central roles. It is worth mentioning that the implementation of these issues is beyond the scope of the present work, and we intend to take this up in our future endeavor.
Moreover, the predicted enhancements in disk luminosity due to the presence of a dense dark matter spike could be degenerate with other astrophysical factors: black hole spin alters the innermost stable circular orbit and thereby the radiative efficiency Shapiro (1974); McKinney and Gammie (2004); Li et al. (2005); magnetic fields can significantly influence angular momentum transport and energy dissipation (e.g., via magneto-rotational instability) Lynden-Bell (1969); Shapiro (1973); Chakrabarti and Mandal (2006); Sarkar et al. (2018); and time variable accretion rates can mimic or obscure dark matter induced signatures Shakura and Sunyaev (1973); Paczyńsky and Wiita (1980); Chakrabarti and Titarchuk (1995). These degeneracies are particularly pertinent in the context of the EHT, which probes the innermost regions of accretion flows, as well as in indirect dark matter detection efforts, where excess emission may be attributed to non-baryonic processes. Future extensions incorporating spin (via Kerr or Kerr-like metrics) and magnetohydrodynamic (MHD) effects are essential to disentangle dark matter contributions from baryonic degeneracies and to more robustly predict observables that can be confronted with high-resolution EHT data.
@@ IX Acknowledgment
Authors thank the anonymous reviewer for useful suggestions and comments that help to improve the quality of the paper. The work of AC is supported by the National Postdoctoral Fellowship of the Science and Engineering Research Board (SERB), ANRF, Govt. of India (File No.: PDF/2023/000550). SC acknowledges support of MATRICS research grant awarded by Science and Engineering Research Board (SERB), ANRF, Govt. of India through grant no. MTR/2022/000318.
@@ Appendix A Keplerian angular momentum
This implies that matter rotating in a Keplerian orbit will stay in equilibrium as the centrifugal force balances the effective gravitational attraction. The variation of the Keplerian angular momentum for the different DM profiles is shown in Fig.˜11. We note that the presence of the DM halo always increases the Keplerian angular momentum from that of Schwarzschild BH. This implies that advection-dominated transonic hot accretion flow in the presence of DM halo can have higher specific angular momentum. Except for highly compact halos, the Einasto distribution shows the highest Keplerian angular momentum at any radius.
The minima of the Keplerian angular momenta correspond to the radius of the marginally stable orbit Abramowicz and Fragile (2013); Rezzolla and Zanotti (2013). We observe that the radius of the marginally stable orbit (ISCO, see Eq.˜7) shifts inwards due to the presence of the DM halo, signalling a weaker inflow. This is consistent with the observation in Heydari-Fard et al. (2025).